# Learning to Generate Object Interactions with Physics-Guided Video Diffusion

### v1

@@ A.2 Additional details
Figure 11: User study interface. We use a telegram bot to display pairwise comparisons of videos and baselines, mixing their order. We collect pairwise user preference on direction, interactions realism, and consistency of objects.
A.3 Additional results
@@
Table 3: Additional ablation studies. We report an additional encoding strategies for our conditioning masks in Table (3(a)), proving that our design choice of using $v_{t}$ is best. Furthermore, we showcase that removing our non-uniform sampling on interactions-rich frames during dropout harms performance (3(b)).
Velocity encoding
@@
In Section 4.1, we mentioned that we sample non-uniformly the frame $f^{*}$ for our second-stage training. In Table 3(b), we present a quantitative evaluation of the effects of this design choice. As visible, sampling the dropout around the frames where collisions occur in the training set helps KineMask to focus on interaction-rich representations.
@@
In Figure 12, we report an interesting failure mode of the concurrent Force Prompting succesfully handled by KineMask. As visible, Force Prompting struggles when the control is applied to object with thin support structures, or with piled objects. This is the result of an architectural decision, since their control mechanism does not employ masks for identifying the object to move. Instead, KineMask correctly generates motion on these edge cases.
KineMask Force Prompting
Figure 12: Motion control comparison. Our mask-based control is robust to ambiguous scenes where objects overlap, while Force Prompting (the second-best method) suffers from hallucinations (right) due to the ambiguous mapping of the control signal to objects in the scene.
@@
Figure 13 shows some failure cases of KineMask, other similar examples can be found in Supplementary Materials. In the first case, we have found that objects that do not have a considerable height tend to ignore others during their motion, thus a collision is not created. The second and third cases, show complex scenarios with many object. This sometimes creates ambiguities, by resulting in object duplication or disappearance. We speculate here that the text prompt may also encourage ambiguity, due to the presence of multiple elements that can be associated to the same textual identifier.
@@ A.4 LLM usage
In this research, we used an LLM, specifically GPT-5, to aid writing. The usage was limited to polishing some sentences and grammatical checks, as well as to identifying inconsistencies in the used terminologies.
Figure 13: Failure Cases

### v2

@@ A.3 Additional results
Figure 14: User study interface. We collect pairwise user preference on direction, interactions realism, and consistency of objects.
Additional Qualitative Comparison
In addition to comparing with our initial set of baselines, we also perform an additional qualitative comparison with the recent PhysGen3D  [13], which uses 3D information, a simulator in the loop and a diffusion model as a rendering tool to generate plausible videos from an initial image. While this brings plausible physics due to the dynamics simulator, this causes multiple problems and limitations, from the complexity of the pipeline, dependency on the 3D recontruction, and the lack of visual realism. As shown in Figure 13, PhysGen3D is vulnerable and dependent on its 3D reconstruction setup, which can cause incorrect object reconstructions, such as creating duplicated objects, like in the first case where the beer is created twice, or in the third example where the reconstruction is incorrect, causing a chain of failures towards synthesizing proper rigid-body interactions. Beyond the synthetic appearance to objects, PhysGen3D and simulation-based methods also are unabile to generate effects like liquid or steam moving as a result of a collision, as shown in the second example where two cups collide. On the other hand, KineMask synthesizes plausible object collisions and effects created from object interactions, showing its understanding of rigid-body dynamics.
Velocity encoding
@@
Table 5: Additional ablation studies. We report an additional encoding strategies for our conditioning masks in Table (5(a)), proving that our design choice of using $v_{t}$ is best. Furthermore, we showcase that removing our non-uniform sampling on interactions-rich frames during dropout harms performance (5(b)).
@@
Section 4.1 mentioned that we sample non-uniformly the frame $f^{*}$ for our second-stage training. In Table 5(b), we present a quantitative evaluation of the effects of this design. As visible, sampling the dropout around the frames where collisions occur in the training set helps KineMask to focus on interaction-rich representations.
@@
In Figure 15, we report an interesting failure mode of drag-based baselines succesfully handled by KineMask. As visible, Force Prompting, TORA and Motion-I2V struggle when the control is applied to object with thin support structures, or with piled objects. This is the result of an architectural decision, since their control mechanism does not employ masks for identifying the object to move with pixel level precision. Instead, KineMask correctly generates motion on these edge cases. This shows how the presence of a dense mask brings a lot of benefits to perform better motion control in ambiguous cases compared to sparse conditioning information.
@@
Figure 17 shows some failure cases of KineMask. In the first and second case, we have found that objects that do not have a considerable height tend to ignore others during their motion, thus a collision is not created. The third and fourth cases, show complex scenarios with many objects. This sometimes creates ambiguities, by resulting in object duplication or disappearance. We speculate here that the text prompt may also encourage ambiguity, due to the presence of multiple elements that can be associated to the same textual identifier.
@@ A.4 LLM usage
In this research, we used GPT-5, to aid writing. The usage was limited to polishing some sentences and grammatical checks, as well as to identifying inconsistencies in the used terminologies.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 15: Motion control comparison. Our mask-based control is robust to ambiguous scenes, while other methods suffer from hallucinations due to the ambiguous control signal.
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@
Figure 17: Failure Cases. We noticed that KineMasks fails to generate collisions with objects that do not have a considerable heigh and in ambiguous setups that contain multiple similar objects.